# Multispectral Pedestrian Detection with Sparsely Annotated Label

### v1

@@ Abstract
Although existing Sparsely Annotated Object Detection (SAOD) approches have made progress in handling sparsely annotated environments in multispectral domain, where only some pedestrians are annotated, they still have the following limitations: (i) they lack considerations for improving the quality of pseudo-labels for missing annotations, and (ii) they rely on fixed ground truth annotations, which leads to learning only a limited range of pedestrian visual appearances in the multispectral domain. To address these issues, we propose a novel framework called Sparsely Annotated Multispectral Pedestrian Detection (SAMPD). For limitation (i), we introduce Multispectral Pedestrian-aware Adaptive Weight (MPAW) and Positive Pseudo-label Enhancement (PPE) module. Utilizing multispectral knowledge, these modules ensure the generation of high-quality pseudo-labels and enable effective learning by increasing weights for high-quality pseudo-labels based on modality characteristics. To address limitation (ii), we propose an Adaptive Pedestrian Retrieval Augmentation (APRA) module, which adaptively incorporates pedestrian patches from ground-truth and dynamically integrates high-quality pseudo-labels with the ground-truth, facilitating a more diverse learning pool of pedestrians. Extensive experimental results demonstrate that our SAMPD significantly enhances performance in sparsely annotated environments within the multispectral domain.
@@ Introduction
When pedestrian annotations are incomplete, it can lead to inconsistencies where pedestrians are misclassified as either foreground or background. To tackle this issue, Sparsely Annotated Object Detection (SAOD) has introduced. Pseudo label methods (Niitani et al. 2019) establish logical connections between object co-occurrence and pseudo-label application. BRL (Zhang et al. 2020) introduces automatic adjustments for areas prone to mislabeling. Co-mining (Wang et al. 2021) jointly traines models using predictions from both original and augmented images. SparseDet (Suri et al. 2023) introduces a self-supervised loss to prevent negative gradient propagation. Calibrated Teacher (Wang et al. 2023a) validates pseudo-label quality using a calibrator that distinguishes positive from negative labels. However, existing methods rely on the pseudo-labels without considering their quality. In contrast, our SAMPD enhances pseudo-label quality by integrating multispectral knowledge and incorporating filtered high-quality pseudo-labels through the Adaptive Pedestrian Retrieval Augmentation (APRA) module.
Figure 2: Overall architecture of our SAMPD. Based on the input image, the augmented image is generated by the APRA module, and the augmented image is used as the actual input for the training phase. $\bm{\oplus}$ denotes the concatenation. During inference phase, only student model is used for detection, and only the original input image (without augmentation) is utilized.
@@
However, some pseudo-labels from the teacher model may be incorrect, capturing only parts of pedestrians or including background elements. These errors can negatively impact the performance of the student model and cause confusion during training. To address this, we propose a multispectral pedestrian-aware adaptive weight (MPAW) module to increase the influence of reliable pseudo-labels and reduces the impact of unreliable ones for each modality. This approach emphasizes high-quality pseudo-labels and reliable modalities. For each modality $k=\{V,T,F\}$, we extract $N$ feature maps from the pseudo-label boxes of the student model $\textbf{f}_{PL}^{k(s)}=\{f_{PL_{i}}^{k(s)}\}_{i=1}^{N}$ and $M$ feature maps from the ground-truth labels $\textbf{f}_{GT}^{k(s)}=\{f_{GT_{i}}^{k(s)}\}_{i=1}^{M}$. We then apply global average pooling (GAP) to obtain latent vectors $\textbf{l}_{PL}^{k(s)}=\{l_{PL_{i}}^{k(s)}\}_{i=1}^{N}$ and $\textbf{l}_{GT}^{k(s)}=\{l_{GT_{i}}^{k(s)}\}_{i=1}^{M}$. We compute the modality weight for $k$ modality, $w^{k}$, as defined by:
@@
Pedestrian Augmentation. To see the effectiveness of our APRA module, we compared it with the recent pedestrian augmentation approach, Robust Teacher (Li et al. 2023). As shown in Table 4, while the Robust Teacher improved performance over the baseline (‘Supervised’), our dynamic approach of APRA module outperformed it. In fact, since Robust Teacher does not account for scene-specific characteristics (e.g., scene brightness, patch sizes, etc.) leads to bounadry bias and visual distortion from augmentation that do not match the lighting conditions of the scene. As a result, it performs worse compared to the static approach (w/o dynamic refinement).
@@ Conclusion
Acknowledgement
This work was supported by the NRF grant funded by the Korea government (MSIT) (No. RS-2023-00252391), and by IITP grant funded by the Korea government (MSIT) (No. RS-2022-00155911: Artificial Intelligence Convergence Innovation Human Resources Development (Kyung Hee University), IITP-2022-II220078: Explainable Logical Reasoning for Medical Knowledge Generation), and by the MSIT (Ministry of Science and ICT), Korea, under the National Program for Excellence in SW (2023-0-00042) supervised by the IITP in 2025, and conducted by CARAI grant funded by DAPA and ADD (UD230017TD).
@@ Acknowledgement
Source Code
The source code of our work will be released as soon as possible after the review process is over.

### v3

@@ Abstract
Although existing Sparsely Annotated Object Detection (SAOD) approaches have made progress in handling sparsely annotated environments in multispectral domain, where only some pedestrians are annotated, they still have the following limitations: (i) they lack considerations for improving the quality of pseudo-labels for missing annotations, and (ii) they rely on fixed ground truth annotations, which leads to learning only a limited range of pedestrian visual appearances in the multispectral domain. To address these issues, we propose a novel framework called Sparsely Annotated Multispectral Pedestrian Detection (SAMPD). For limitation (i), we introduce Multispectral Pedestrian-aware Adaptive Weight (MPAW) and Positive Pseudo-label Enhancement (PPE) module. Utilizing multispectral knowledge, these modules ensure the generation of high-quality pseudo-labels and enable effective learning by increasing weights for high-quality pseudo-labels based on modality characteristics. To address limitation (ii), we propose an Adaptive Pedestrian Retrieval Augmentation (APRA) module, which adaptively incorporates pedestrian patches from ground-truth and dynamically integrates high-quality pseudo-labels with the ground-truth, facilitating a more diverse learning pool of pedestrians. Extensive experimental results demonstrate that our SAMPD significantly enhances performance in sparsely annotated environments within the multispectral domain. The code is available at https://github.com/VisualAIKHU/SAMPD.
@@ Introduction
When pedestrian annotations are incomplete, it can lead to inconsistencies where pedestrians are misclassified as either foreground or background. To tackle this issue, Sparsely Annotated Object Detection (SAOD) has introduced. Pseudo label methods (Niitani et al. 2019) establish logical connections between object co-occurrence and pseudo-label application. BRL (Zhang et al. 2020) introduces automatic adjustments for areas prone to mislabeling. Co-mining (Wang et al. 2021) jointly trains models using predictions from both original and augmented images. SparseDet (Suri et al. 2023) introduces a self-supervised loss to prevent negative gradient propagation. Calibrated Teacher (Wang et al. 2023a) validates pseudo-label quality using a calibrator that distinguishes positive from negative labels. However, existing methods rely on the pseudo-labels without considering their quality. In contrast, our SAMPD enhances pseudo-label quality by integrating multispectral knowledge and incorporating filtered high-quality pseudo-labels through the Adaptive Pedestrian Retrieval Augmentation (APRA) module.
Figure 2: Overall architecture of our SAMPD. Given the multispectral input image, the APRA module generates an augmented image, which is then used as the input for the training phase. $\bm{\oplus}$ denotes the concatenation. During inference phase, only student model is used for detection, and only the original input image (without augmentation) is utilized.
@@
However, some pseudo-labels from the teacher model may be incorrect, capturing only parts of pedestrians or including background elements. These errors can negatively impact the performance of the student model and cause confusion during training. To address this, we propose a multispectral pedestrian-aware adaptive weight (MPAW) module to increase the influence of reliable pseudo-labels and reduce the impact of unreliable ones for each modality. This approach emphasizes high-quality pseudo-labels and reliable modalities. For each modality $k=\{V,T,F\}$, we extract $N$ feature maps from the pseudo-label boxes of the student model $\textbf{f}_{PL}^{k(s)}=\{f_{PL_{i}}^{k(s)}\}_{i=1}^{N}$ and $M$ feature maps from the ground-truth labels $\textbf{f}_{GT}^{k(s)}=\{f_{GT_{i}}^{k(s)}\}_{i=1}^{M}$. We then apply global average pooling (GAP) to obtain latent vectors $\textbf{l}_{PL}^{k(s)}=\{l_{PL_{i}}^{k(s)}\}_{i=1}^{N}$ and $\textbf{l}_{GT}^{k(s)}=\{l_{GT_{i}}^{k(s)}\}_{i=1}^{M}$. We compute the modality weight for $k$ modality, $w^{k}$, as defined by:
@@
Pedestrian Augmentation. To see the effectiveness of our APRA module, we compared it with the recent pedestrian augmentation approach, Robust Teacher (Li et al. 2023). As shown in Table 4, while the Robust Teacher improved performance over the baseline (‘Supervised’), our dynamic approach of APRA module outperformed it. In fact, since Robust Teacher does not account for scene-specific characteristics (e.g., scene brightness, patch sizes, etc.) leads to boundary bias and visual distortion from augmentation that do not match the lighting conditions of the scene. As a result, it performs worse compared to the static approach (w/o dynamic refinement).
@@ Conclusion
Acknowledgements
This work was supported by the NRF grant funded by the Korea government (MSIT) (No. RS-2023-00252391), and by IITP grant funded by the Korea government (MSIT) (No. RS-2022-00155911: Artificial Intelligence Convergence Innovation Human Resources Development (Kyung Hee University), IITP-2022-II220078: Explainable Logical Reasoning for Medical Knowledge Generation, No. RS-2024-00509257: Global AI Frontier Lab), and by the MSIT (Ministry of Science and ICT), Korea, under the National Program for Excellence in SW (2023-0-00042) supervised by the IITP in 2025, and conducted by CARAI grant funded by DAPA and ADD (UD230017TD).